OpenStudy (anonymous):

state the original pythagorean identity and then show the steps to transform the identity into two additional pythagorean identities

OpenStudy (mertsj):

\[\sin ^2x+\cos ^2x=1\]

OpenStudy (anonymous):

that was the original correct?

OpenStudy (mertsj):

\[\frac{\sin ^2x}{\sin ^2x}+\frac{\cos ^2x}{\sin ^2x}=\frac{1}{\sin ^2x}\] \[1+\cot ^2x=\csc ^2x\]

OpenStudy (mertsj):

\[\frac{\sin ^2x}{\cos ^2x}+\frac{\cos ^2x}{\cos ^2x}=\frac{1}{\cos ^2x}\] \[\tan ^2x+1=\sec ^2x\]